# HiPhO: How Far Are (M)LLMs from Humans in the Latest High School Physics Olympiad Benchmark?

### v2

@@ Abstract
Recently, the physical capabilities of (M)LLMs have garnered increasing attention. However, existing benchmarks for physics suffer from two major gaps: they neither provide systematic and up-to-date coverage of real-world physics competitions such as physics Olympiads, nor enable direct performance comparison with humans. To bridge these gaps, we present HiPhO, the first benchmark dedicated to high school physics Olympiads with human-aligned evaluation. Specifically, HiPhO highlights three key innovations. (1) Comprehensive Data: It compiles 13 latest Olympiad exams from 2024–2025, spanning both international and regional competitions, and covering mixed modalities that encompass problems spanning text-only to diagram-based. (2) Professional Evaluation: We adopt official marking schemes to perform fine-grained grading at both the answer and step level, fully aligned with human examiners to ensure high-quality and domain-specific evaluation. (3) Comparison with Human Contestants: We assign gold, silver, and bronze medals to models based on official medal thresholds, thereby enabling direct comparison between (M)LLMs and human contestants. Our large-scale evaluation of 30 state-of-the-art (M)LLMs shows that: across 13 exams, open-source MLLMs mostly remain at or below the bronze level; open-source LLMs show promising progress with occasional golds; closed-source reasoning MLLMs can achieve 6 to 12 gold medals; and most models still have a significant gap from full marks. These results highlight a substantial performance gap between open-source models and top students, the strong physical reasoning capabilities of closed-source reasoning models, and the fact that there is still significant room for improvement. HiPhO, as a rigorous, human-aligned, and Olympiad-focused benchmark for advancing multimodal physical reasoning, is open-source and available at https://github.com/SciYu/HiPhO.
@@ 1 Introduction
In our large-scale evaluation of 30 state-of-the-art (M)LLMs, we observe a clear performance hierarchy. Closed-source reasoning MLLMs dominate the medal table, winning 6–12 gold medals in 13 Olympiads. However, even the strongest models, such as Gemini-2.5-Pro and GPT-5, still fall short of the very best human contestants, especially in challenging exams like IPhO and EuPhO. In contrast, open-source chat MLLMs failed to secure any gold medals, with most scoring only at or below the bronze threshold. More encouragingly, several open-source reasoning (M)LLMs, including Intern-S1 and DeepSeek-R1, each achieved 4–5 gold medals, particularly in relatively easier exams such as F=MA. Taken together, these results reveal the strong but still non-parity performance of closed-source models, the evident limitations of open-source chat MLLMs, and the promising trajectory of open-source reasoning (M)LLMs in advancing physics problem solving.
@@
$\bullet$ Human-level Comparison. Compared to human contestants, closed-source reasoning MLLMs reach gold in 6–12 exams, while open-source MLLMs stay mostly at bronze; some open-source LLMs show stronger reasoning with occasional golds, yet all remain far from the very top students.
@@ 4 Evaluation Method
GPT-OSS-120B 15.2 21.0 20.0 9.1 18.6 20.9 22.2 37.4 41.8 57.1 59.7 17.8 17.6  5 7 1
DeepSeek-R1 18.1 24.6 22.3 10.5 21.3 24.0 20.0 22.4 20.3 48.5 63.8 18.3 18.5  5 6 2
Qwen3-235B-A22B 17.5 23.8 23.0 9.2 20.6 25.4 30.3 26.0 18.9 48.9 40.4 18.4 17.4  5 5 3
Kimi-K2-Instruct 15.1 19.2 20.3 10.1 15.2 22.7 23.6 35.9 41.8 65.9 58.9 16.0 18.2  4 9 0
Qwen3-32B 15.4 19.3 21.3 9.8 20.0 25.0 22.6 26.8 17.9 50.1 58.7 18.9 16.6  3 8 2
Qwen3-30B-A3B 12.9 15.4 20.4 9.8 15.0 21.9 20.2 22.8 17.8 40.3 44.6 15.5 15.0  2 9 2
DeepSeek-V3 13.5 16.4 18.8 7.0 15.1 18.6 16.8 23.5 16.9 36.7 44.8 14.1 15.6  1 7 5
Qwen3-8B 10.2 12.6 9.9 7.1 11.1 19.0 9.7 21.2 11.5 15.0 22.5 10.8 10.0  0 1 8
@@ 5.1 Main Results
What are the most powerful (M)LLMs in physics Olympiads? The medal table in Table 3 ranks models just like an Olympiad. The top five positions are occupied by Gemini-2.5-Pro, Gemini-2.5-Flash, GPT-5, o3, and Grok-4. However, in the challenging IPhO-2025, only three of them reached the gold threshold, with the top two separated by merely 0.4 points. This narrow margin highlights the intense competition at the frontier of closed-source model performance.
@@
Can open-source models catch up with closed-source champions? While closed-source reasoning models still dominate, recent progress in the open-source community is noteworthy. Intern-S1 distinguished itself as the only open-source MLLM to win gold, obtaining four medals. Even more impressively, open-source LLMs including GPT-OSS-120B, DeepSeek-R1, and Qwen3-235B-A22B all achieved five gold medals. These advances reveal both the enduring gap with closed-source leaders and the emerging potential of open-source research to narrow it.
@@ 5.2 Can SOTA (M)LLMs Compete with Human Medalists in Physics Olympiads?
SOTA open-source (M)LLMs usually reach bronze to silver levels, but still fall short of human gold medalists in most Olympiads. As shown in Fig. 5, open-source models delivered respectable results in several contests, often scoring near the median of human medalists. For instance, in PanPhO-2025 Intern-S1 scored above 40 points, surpassing the median performance of human medalists, yet still below the gold threshold. By contrast, in more challenging Olympiads such as EuPhO-2025, which featured optics problems with variable figures, open-source MLLMs dropped back to the bronze level. These results illustrate both the steady progress of open-source models and the considerable gap that remains between them and the top-performing human contestants.
@@ 5.4 Which Physics Fields Challenge MLLMs the Most?
Physics Olympiad problems span five fields: Mechanics (Mech.), Electromagnetism (Elec.), Thermodynamics (Ther.), Optics (Opt.), and Modern Physics (Mode.). We report mean normalized scores (Eq. 1) for each field in Table 4, leading to three key observations. (1) Optics is the most challenging field, with all models scoring below 55%. Its difficulty arises from two aspects: geometrical optics requires diagram interpretation, while wave optics demands precise symbolic derivations—both remain weak points for current MLLMs. (2) Modern Physics tends to yield higher scores than Mechanics, likely because the problems are less visually intensive and the concepts tested do not extend to university-level depth. (3) Mechanics and Electromagnetism show a clear source gap. Closed-source models average above 70%. Most open-source MLLMs remain below 45%, with one exception: Intern-S1 narrows the gap, reaching 58–60% in Mech./Elec. Overall, optics is the most challenging, while closed-source models show steady gains.
@@ 6 Conclusion and Future Work
In this work, we presented HiPhO, the first benchmark designed to systematically evaluate MLLMs against top-performing students in high school physics Olympiads. Spanning 13 Olympiad exams from 2024–2025, HiPhO covers both international and regional contests. The problems include mixed modalities, spanning text-only to diagram-based, with fine-grained categorization. HiPhO is the first to combine answer-level and step-level grading based on official marking schemes, enabling rigorous, human-aligned evaluation. We compare model exam scores against human contestants and observe significant performance gaps. SOTA closed-source reasoning models achieve gold-level performance on most exams but still fall short of the very best human contestants. Most open-source models remain at the bronze level or below, with occasional golds on easier contests. These findings highlight both the progress of MLLMs in physical reasoning and the substantial gap that remains before achieving human-level mastery.
@@ C.1 Experimental Setup
Source Developer Reasoning Models Chat (Non-Reasoning) Models
@@
Google Gemini-2.5-Pro, Gemini-2.5-Flash

### v3

@@ Abstract
Recently, the physical capabilities of (M)LLMs have garnered increasing attention. However, existing benchmarks for physics suffer from two major gaps: they neither provide systematic and up-to-date coverage of real-world physics competitions such as physics Olympiads, nor enable direct performance comparison with humans. To bridge these gaps, we present HiPhO, the first benchmark dedicated to high school physics Olympiads with human-aligned evaluation. Specifically, HiPhO highlights three key innovations. (1) Comprehensive Data: It compiles 13 latest Olympiad exams from 2024–2025, spanning both international and regional competitions, and covering mixed modalities that encompass problems spanning text-only to diagram-based. (2) Professional Evaluation: We adopt official marking schemes to perform fine-grained grading at both the answer and step level, fully aligned with human examiners to ensure high-quality and domain-specific evaluation. (3) Comparison with Human Contestants: We assign gold, silver, and bronze medals to models based on official medal thresholds, thereby enabling direct comparison between (M)LLMs and human contestants. Our large-scale evaluation of 30 state-of-the-art (M)LLMs shows that: across 13 exams, open-source MLLMs mostly remain at or below the bronze level; open-source LLMs show promising progress with multiple golds; closed-source reasoning MLLMs can achieve 6 to 12 gold medals; and most models still have a significant gap from full marks. These results highlight the performance gap between open-source models and top students, the strong reasoning abilities of closed-source models, and the remaining room for improvement. HiPhO, a human-aligned Olympiad benchmark for multimodal physical reasoning, is open-source at https://github.com/SciYu/HiPhO with a public leaderboard at https://phyarena.github.io/.
@@ 1 Introduction
In our large-scale evaluation of 30 state-of-the-art (M)LLMs, we observe a clear performance hierarchy. Closed-source reasoning MLLMs dominate the medal table, winning 6–12 gold medals in 13 Olympiads. However, even the strongest models, such as Gemini-2.5-Pro and GPT-5, still fall short of the very best human contestants, especially in challenging exams like IPhO and EuPhO. In contrast, open-source chat MLLMs failed to secure any gold medals, with most scoring only at or below the bronze threshold. More encouragingly, several open-source reasoning (M)LLMs, including Intern-S1 and DeepSeek-R1, each achieved 4–6 gold medals, particularly in relatively easier exams such as F=MA. Taken together, these results reveal the strong but still non-parity performance of closed-source models, the evident limitations of open-source chat MLLMs, and the promising trajectory of open-source reasoning (M)LLMs in advancing physics problem solving.
@@
$\bullet$ Human-level Comparison. Compared to human contestants, closed-source reasoning MLLMs reach gold in 6–12 exams, while open-source MLLMs stay mostly at bronze; some open-source LLMs show stronger reasoning with multiple golds, yet all remain far from the very top students.
@@ 4 Evaluation Method
DeepSeek-R1 18.5 24.6 25.4 10.8 21.4 26.3 20.5 22.4 20.3 48.5 63.8 18.3 18.5  6 5 2
Qwen3-32B 15.7 19.3 23.9 9.8 21.2 28.9 24.1 26.8 17.9 50.1 58.7 18.9 16.6  6 5 2
Qwen3-235B-A22B 17.8 23.8 26.0 9.2 21.5 28.4 31.1 26.0 18.9 48.9 40.4 18.4 17.4  6 4 3
Kimi-K2-Instruct 16.5 19.8 24.2 11.0 16.9 26.5 26.2 35.9 41.8 65.9 58.9 16.0 18.2  5 8 0
GPT-OSS-120B 16.9 21.4 22.8 9.1 19.9 26.0 25.8 37.4 41.8 57.1 59.7 17.8 17.6  5 7 1
Qwen3-30B-A3B 13.6 15.4 22.7 9.8 16.5 24.7 21.7 22.8 17.8 40.3 44.6 15.5 15.0  2 9 2
DeepSeek-V3 13.6 16.4 22.1 7.1 17.2 21.1 17.3 23.5 16.9 36.7 44.8 14.1 15.6  1 8 4
Qwen3-8B 10.6 12.7 11.5 7.1 11.9 20.1 17.3 21.2 11.5 15.0 22.5 10.8 10.0  0 2 8
@@ 5.1 Main Results
What are the most powerful (M)LLMs in physics Olympiads? The medal table in Table 3 ranks models just like an Olympiad. The top five positions are occupied by Gemini-2.5-Pro, Gemini-2.5-Flash-Thinking, GPT-5, o3, and Grok-4. However, in the challenging IPhO-2025, only three of them reached the gold threshold, with the top two separated by merely 0.4 points. This narrow margin highlights the intense competition at the frontier of closed-source model performance.
@@
Can open-source models catch up with closed-source champions? While closed-source reasoning models still dominate, recent progress in the open-source community is noteworthy. Intern-S1 distinguished itself as the only open-source MLLM to win gold, obtaining four medals. Even more impressively, open-source LLMs including DeepSeek-R1, Qwen3-32B, and Qwen3-235B-A22B all achieved six gold medals. These advances reveal both the enduring gap with closed-source leaders and the emerging potential of open-source research to narrow it.
@@ 5.2 Can SOTA (M)LLMs Compete with Human Medalists in Physics Olympiads?
SOTA open-source MLLMs usually reach bronze to silver levels, but still fall short of human gold medalists in most Olympiads. As shown in Fig. 5, open-source models delivered respectable results in several contests, often scoring near the median of human medalists. For instance, in PanPhO-2025 Intern-S1 scored above 40 points, surpassing the median performance of human medalists, yet still below the gold threshold. By contrast, in more challenging Olympiads such as EuPhO-2025, which featured optics problems with variable figures, open-source MLLMs dropped back to the bronze level. These results illustrate both the steady progress of open-source models and the considerable gap that remains between them and the top-performing human contestants.
@@ 5.4 Which Physics Fields Challenge MLLMs the Most?
Physics Olympiad problems span five fields: Mechanics (Mech.), Electromagnetism (Elec.), Thermodynamics (Ther.), Optics (Opt.), and Modern Physics (Mode.). We report mean normalized scores (Eq. 1) for each field in Table 4, leading to three key observations. (1) Optics is the most challenging field, with all models scoring below 55%. Its difficulty arises from two aspects: geometrical optics requires diagram interpretation, while wave optics demands precise symbolic derivations—both remain weak points for current MLLMs. (2) Modern Physics tends to yield higher scores than Mechanics, likely because the problems are less visually intensive and the concepts tested do not extend to university-level depth. (3) Mechanics and Electromagnetism show a clear source gap. Closed-source models average above 70%. Most open-source MLLMs remain below 45%, with one exception: Intern-S1 narrows the gap, reaching 63–67% in Mech./Elec. Overall, optics is the most challenging, while closed-source models show steady gains.
@@ 6 Conclusion and Future Work
In this work, we presented HiPhO, the first benchmark designed to systematically evaluate (M)LLMs against top-performing students in high school physics Olympiads. Spanning 13 Olympiad exams from 2024–2025, HiPhO covers both international and regional contests. The problems include mixed modalities, spanning text-only to diagram-based, with fine-grained categorization. HiPhO is the first to combine answer-level and step-level grading based on official marking schemes, enabling rigorous, human-aligned evaluation. We compare model exam scores against human contestants and observe significant performance gaps. SOTA closed-source reasoning MLLMs achieve gold-level performance on most exams but still fall short of the very best human contestants. Most open-source MLLMs remain at the bronze level or below, while open-source LLMs obtain multiple golds. These findings highlight both the progress of (M)LLMs in physical reasoning and the substantial gap that remains before achieving human-level mastery.
@@ C.1 Experimental Setup
Source Developer Reasoning Models Chat Models
@@
Google Gemini-2.5-Pro, Gemini-2.5-Flash-Thinking